### v1

@@ Appendix D Simulation results
7.1 $h=2$ $\mathscr{T}_{n}$ 0.088 0.200 0.248 0.094 0.156 0.212 0.090 0.190 0.266
$\mathscr{T}_{n}^{*}$ 0.039 0.059 0.108 0.049 0.088 0.127 0.020 0.039 0.069
$h=4$ $\mathscr{T}_{n}$ 0.060 0.150 0.206 0.084 0.170 0.218 0.086 0.172 0.204
$\mathscr{T}_{n}^{*}$ 0.010 0.039 0.039 0.029 0.049 0.069 0.049 0.127 0.157
$h=8$ $\mathscr{T}_{n}$ 0.062 0.134 0.194 0.062 0.136 0.174 0.068 0.144 0.202
$\mathscr{T}_{n}^{*}$ 0.020 0.078 0.108 0.049 0.078 0.098 0.020 0.127 0.137
7.1 $h=2$ $\mathscr{T}_{n}$ 0.102 0.142 0.182 0.054 0.096 0.152 0.052 0.106 0.136
$\mathscr{T}_{n}^{*}$ 0.029 0.059 0.118 0.059 0.108 0.127 0.020 0.078 0.176
$h=4$ $\mathscr{T}_{n}$ 0.098 0.158 0.212 0.042 0.090 0.136 0.036 0.086 0.114
$\mathscr{T}_{n}^{*}$ 0.020 0.069 0.118 0.069 0.118 0.137 0.029 0.108 0.147
$h=8$ $\mathscr{T}_{n}$ 0.120 0.210 0.282 0.066 0.122 0.162 0.034 0.088 0.144
$\mathscr{T}_{n}^{*}$ 0.010 0.059 0.088 0.049 0.127 0.157 0.029 0.069 0.167
7.1 $h=2$ $\mathscr{T}_{n}$ 0.068 0.118 0.142 0.024 0.054 0.084 0.028 0.078 0.092
$\mathscr{T}_{n}^{*}$ 0.039 0.059 0.078 0.000 0.020 0.088 0.069 0.069 0.108
$h=4$ $\mathscr{T}_{n}$ 0.054 0.106 0.164 0.024 0.062 0.098 0.048 0.076 0.108
$\mathscr{T}_{n}^{*}$ 0.039 0.069 0.069 0.020 0.059 0.088 0.039 0.098 0.127
$h=8$ $\mathscr{T}_{n}$ 0.062 0.138 0.174 0.028 0.058 0.092 0.032 0.084 0.132
$\mathscr{T}_{n}^{*}$ 0.039 0.059 0.078 0.029 0.069 0.098 0.039 0.078 0.127
trigonometric series $n=100$ $n=300$ $n=500$
@@
7.1 $h=2$ $\mathscr{T}_{n}$ 0.082 0.160 0.212 0.106 0.198 0.246 0.084 0.176 0.232
$\mathscr{T}_{n}^{*}$ 0.029 0.069 0.098 0.078 0.108 0.127 0.020 0.088 0.118
$h=4$ $\mathscr{T}_{n}$ 0.072 0.146 0.208 0.076 0.160 0.204 0.086 0.150 0.208
$\mathscr{T}_{n}^{*}$ 0.029 0.118 0.118 0.029 0.039 0.078 0.029 0.069 0.098
$h=8$ $\mathscr{T}_{n}$ 0.086 0.144 0.194 0.076 0.154 0.204 0.084 0.152 0.184
$\mathscr{T}_{n}^{*}$ 0.039 0.108 0.137 0.029 0.078 0.098 0.049 0.108 0.118
7.1 $h=2$ $\mathscr{T}_{n}$ 0.086 0.134 0.180 0.056 0.112 0.140 0.056 0.108 0.144
$\mathscr{T}_{n}^{*}$ 0.039 0.078 0.118 0.039 0.108 0.137 0.029 0.049 0.147
$h=4$ $\mathscr{T}_{n}$ 0.102 0.188 0.240 0.040 0.100 0.140 0.052 0.092 0.138
$\mathscr{T}_{n}^{*}$ 0.020 0.059 0.069 0.078 0.137 0.167 0.049 0.108 0.147
$h=8$ $\mathscr{T}_{n}$ 0.112 0.192 0.270 0.052 0.106 0.160 0.032 0.076 0.132
$\mathscr{T}_{n}^{*}$ 0.010 0.029 0.049 0.029 0.167 0.176 0.049 0.118 0.167
7.1 $h=2$ $\mathscr{T}_{n}$ 0.046 0.096 0.120 0.026 0.064 0.098 0.026 0.072 0.102
$\mathscr{T}_{n}^{*}$ 0.039 0.059 0.069 0.020 0.059 0.078 0.059 0.098 0.127
$h=4$ $\mathscr{T}_{n}$ 0.040 0.104 0.150 0.022 0.066 0.104 0.034 0.068 0.118
$\mathscr{T}_{n}^{*}$ 0.020 0.078 0.118 0.049 0.088 0.088 0.059 0.088 0.127
$h=8$ $\mathscr{T}_{n}$ 0.056 0.132 0.158 0.022 0.064 0.096 0.024 0.076 0.130
$\mathscr{T}_{n}^{*}$ 0.029 0.069 0.078 0.029 0.078 0.127 0.049 0.088 0.108
Table 16: Empirical sizes for $\mathbb{W}_{3}$ with $d_{\lambda}=1$, rejection probabilities $\mathscr{T}_{n}$ at 1%, 5% and 10% level.
polynomial series $n=100$ $n=300$ $n=500$
@@
$r=2$ 7.1 $h=2$ 0.028 0.092 0.160 0.030 0.096 0.116 0.012 0.060 0.106
$h=4$ 0.020 0.060 0.084 0.024 0.052 0.084 0.028 0.056 0.072
$h=8$ 0.008 0.032 0.048 0.020 0.032 0.048 0.024 0.040 0.072
7.1 $h=2$ 0.044 0.096 0.152 0.044 0.096 0.120 0.014 0.058 0.116
$h=4$ 0.040 0.080 0.108 0.032 0.104 0.116 0.012 0.040 0.088
$h=8$ 0.024 0.044 0.064 0.008 0.016 0.028 0.032 0.060 0.088
7.1 $h=2$ 0.036 0.100 0.112 0.024 0.076 0.104 0.048 0.092 0.132
$h=4$ 0.016 0.052 0.092 0.028 0.048 0.068 0.028 0.068 0.104
$h=8$ 0.012 0.036 0.048 0.032 0.056 0.068 0.004 0.028 0.052
$r=4$ 7.1 $h=2$ 0.032 0.116 0.156 0.040 0.084 0.108 0.018 0.050 0.102
$h=4$ 0.032 0.060 0.088 0.032 0.076 0.088 0.016 0.028 0.080
$h=8$ 0.012 0.032 0.048 0.012 0.032 0.048 0.004 0.032 0.098
7.1 $h=2$ 0.032 0.104 0.136 0.032 0.076 0.092 0.014 0.062 0.104
$h=4$ 0.040 0.100 0.108 0.052 0.092 0.116 0.008 0.028 0.076
$h=8$ 0.024 0.056 0.064 0.008 0.040 0.056 0.024 0.056 0.110
7.1 $h=2$ 0.060 0.096 0.136 0.028 0.080 0.132 0.014 0.048 0.102
$h=4$ 0.036 0.064 0.088 0.008 0.044 0.056 0.024 0.068 0.112
$h=8$ 0.020 0.044 0.060 0.020 0.056 0.080 0.016 0.048 0.094
Table 17: Empirical sizes for $\mathbb{W}_{4}$, rejection probabilities $\mathscr{T}_{n}$ and $\mathscr{T}_{n}^{*}$ at 1%, 5% and 10% level. Testing $\delta(z)$.
polynomial series $n=100$ $n=300$ $n=500$
@@
7.1 $h=2$ $\mathscr{T}_{n}$ 0.036 0.088 0.128 0.064 0.104 0.156 0.044 0.112 0.152
$\mathscr{T}_{n}^{*}$ 0.008 0.028 0.064 0.004 0.022 0.074 0.012 0.036 0.082
$h=4$ $\mathscr{T}_{n}$ 0.044 0.080 0.116 0.016 0.048 0.080 0.040 0.076 0.096
$\mathscr{T}_{n}^{*}$ 0.000 0.024 0.076 0.000 0.040 0.072 0.012 0.036 0.084
$h=8$ $\mathscr{T}_{n}$ 0.008 0.024 0.040 0.004 0.008 0.012 0.004 0.016 0.032
$\mathscr{T}_{n}^{*}$ 0.020 0.064 0.072 0.008 0.020 0.076 0.024 0.052 0.090
7.1 $h=2$ $\mathscr{T}_{n}$ 0.048 0.124 0.160 0.032 0.060 0.108 0.060 0.136 0.172
$\mathscr{T}_{n}^{*}$ 0.000 0.016 0.024 0.004 0.012 0.016 0.016 0.028 0.060
$h=4$ $\mathscr{T}_{n}$ 0.060 0.112 0.132 0.028 0.068 0.104 0.044 0.096 0.128
$\mathscr{T}_{n}^{*}$ 0.000 0.012 0.036 0.000 0.000 0.020 0.008 0.012 0.024
$h=8$ $\mathscr{T}_{n}$ 0.016 0.044 0.056 0.008 0.032 0.056 0.012 0.040 0.068
$\mathscr{T}_{n}^{*}$ 0.000 0.040 0.076 0.004 0.038 0.072 0.006 0.044 0.072
7.1 $h=2$ $\mathscr{T}_{n}$ 0.036 0.104 0.140 0.052 0.116 0.168 0.028 0.104 0.156
$\mathscr{T}_{n}^{*}$ 0.000 0.016 0.068 0.000 0.060 0.076 0.000 0.068 0.094
$h=4$ $\mathscr{T}_{n}$ 0.004 0.016 0.024 0.000 0.004 0.012 0.044 0.068 0.108
$\mathscr{T}_{n}^{*}$ 0.024 0.072 0.100 0.016 0.044 0.080 0.012 0.040 0.088
$h=8$ $\mathscr{T}_{n}$ 0.000 0.028 0.048 0.012 0.032 0.068 0.016 0.036 0.068
$\mathscr{T}_{n}^{*}$ 0.006 0.032 0.084 0.000 0.034 0.078 0.004 0.040 0.076
trigonometric series $n=100$ $n=300$ $n=500$
@@
7.1 $h=2$ $\mathscr{T}_{n}$ 0.052 0.104 0.144 0.044 0.092 0.116 0.048 0.116 0.160
$\mathscr{T}_{n}^{*}$ 0.004 0.022 0.068 0.016 0.020 0.076 0.004 0.034 0.088
$h=4$ $\mathscr{T}_{n}$ 0.012 0.044 0.076 0.032 0.076 0.116 0.052 0.116 0.132
$\mathscr{T}_{n}^{*}$ 0.004 0.024 0.074 0.004 0.030 0.080 0.004 0.038 0.078
$h=8$ $\mathscr{T}_{n}$ 0.000 0.004 0.004 0.000 0.004 0.008 0.008 0.020 0.032
$\mathscr{T}_{n}^{*}$ 0.016 0.036 0.078 0.008 0.020 0.080 0.016 0.028 0.082
7.1 $h=2$ $\mathscr{T}_{n}$ 0.052 0.108 0.132 0.032 0.084 0.108 0.040 0.112 0.148
$\mathscr{T}_{n}^{*}$ 0.000 0.008 0.024 0.000 0.000 0.012 0.008 0.024 0.056
$h=4$ $\mathscr{T}_{n}$ 0.048 0.076 0.116 0.044 0.072 0.104 0.036 0.088 0.128
$\mathscr{T}_{n}^{*}$ 0.008 0.012 0.028 0.000 0.004 0.012 0.000 0.016 0.044
$h=8$ $\mathscr{T}_{n}$ 0.012 0.060 0.072 0.012 0.024 0.040 0.016 0.040 0.116
$\mathscr{T}_{n}^{*}$ 0.000 0.040 0.072 0.004 0.038 0.068 0.008 0.040 0.092
7.1 $h=2$ $\mathscr{T}_{n}$ 0.056 0.128 0.168 0.044 0.084 0.128 0.040 0.108 0.152
$\mathscr{T}_{n}^{*}$ 0.004 0.030 0.064 0.000 0.044 0.088 0.000 0.046 0.094
$h=4$ $\mathscr{T}_{n}$ 0.016 0.024 0.040 0.008 0.056 0.080 0.044 0.104 0.136
$\mathscr{T}_{n}^{*}$ 0.044 0.088 0.100 0.012 0.030 0.068 0.008 0.040 0.094
$h=8$ $\mathscr{T}_{n}$ 0.032 0.044 0.060 0.008 0.036 0.052 0.020 0.040 0.068
$\mathscr{T}_{n}^{*}$ 0.012 0.036 0.082 0.004 0.028 0.062 0.006 0.032 0.068
Table 18: Empirical sizes for $\mathbb{W}_{4}$, rejection probabilities $\mathscr{T}_{n}$ and $\mathscr{T}_{n}^{*}$ at 1%, 5% and 10% level. Testing $\lambda(z)$.
polynomial series $n=100$ $n=300$ $n=500$
1% 5% 10% 1% 5% 10% 1% 5% 10%
7.1 $h=2$ $\mathscr{T}_{n}$ 0.056 0.136 0.168 0.048 0.096 0.140 0.060 0.124 0.156
$\mathscr{T}_{n}^{*}$ 0.024 0.048 0.088 0.020 0.048 0.084 0.004 0.036 0.082
$h=4$ $\mathscr{T}_{n}$ 0.040 0.068 0.096 0.040 0.080 0.100 0.032 0.088 0.132
$\mathscr{T}_{n}^{*}$ 0.016 0.032 0.094 0.012 0.032 0.076 0.004 0.052 0.092
$h=8$ $\mathscr{T}_{n}$ 0.020 0.056 0.100 0.024 0.068 0.084 0.040 0.096 0.116
$\mathscr{T}_{n}^{*}$ 0.004 0.036 0.088 0.012 0.036 0.068 0.020 0.036 0.100
7.1 $h=2$ $\mathscr{T}_{n}$ 0.036 0.076 0.108 0.036 0.104 0.136 0.072 0.124 0.168
$\mathscr{T}_{n}^{*}$ 0.016 0.032 0.052 0.012 0.032 0.076 0.020 0.052 0.104
$h=4$ $\mathscr{T}_{n}$ 0.032 0.108 0.152 0.028 0.076 0.120 0.060 0.108 0.136
$\mathscr{T}_{n}^{*}$ 0.016 0.028 0.068 0.008 0.020 0.068 0.020 0.040 0.092
$h=8$ $\mathscr{T}_{n}$ 0.044 0.088 0.100 0.020 0.040 0.106 0.012 0.040 0.096
$\mathscr{T}_{n}^{*}$ 0.020 0.060 0.080 0.012 0.028 0.036 0.012 0.028 0.040
7.1 $h=2$ $\mathscr{T}_{n}$ 0.052 0.136 0.192 0.048 0.116 0.148 0.072 0.184 0.216
$\mathscr{T}_{n}^{*}$ 0.008 0.048 0.080 0.012 0.040 0.068 0.020 0.064 0.108
$h=4$ $\mathscr{T}_{n}$ 0.028 0.064 0.108 0.032 0.076 0.092 0.072 0.144 0.204
$\mathscr{T}_{n}^{*}$ 0.008 0.032 0.072 0.008 0.028 0.056 0.024 0.056 0.128
$h=8$ $\mathscr{T}_{n}$ 0.036 0.072 0.076 0.020 0.076 0.112 0.028 0.064 0.080
$\mathscr{T}_{n}^{*}$ 0.012 0.044 0.088 0.004 0.036 0.076 0.016 0.040 0.086
trigonometric series $n=100$ $n=300$ $n=500$
1% 5% 10% 1% 5% 10% 1% 5% 10%
7.1 $h=2$ $\mathscr{T}_{n}$ 0.040 0.112 0.152 0.048 0.100 0.124 0.044 0.108 0.132
$\mathscr{T}_{n}^{*}$ 0.012 0.036 0.088 0.016 0.040 0.094 0.008 0.058 0.082
$h=4$ $\mathscr{T}_{n}$ 0.048 0.096 0.108 0.012 0.072 0.108 0.056 0.116 0.136
$\mathscr{T}_{n}^{*}$ 0.016 0.044 0.088 0.004 0.038 0.082 0.012 0.048 0.088
$h=8$ $\mathscr{T}_{n}$ 0.032 0.056 0.080 0.028 0.076 0.096 0.028 0.064 0.072
$\mathscr{T}_{n}^{*}$ 0.024 0.036 0.072 0.008 0.036 0.080 0.024 0.036 0.088
7.1 $h=2$ $\mathscr{T}_{n}$ 0.036 0.092 0.136 0.064 0.112 0.172 0.056 0.152 0.200
$\mathscr{T}_{n}^{*}$ 0.008 0.044 0.084 0.004 0.028 0.080 0.016 0.056 0.116
$h=4$ $\mathscr{T}_{n}$ 0.036 0.076 0.104 0.020 0.060 0.084 0.052 0.096 0.112
$\mathscr{T}_{n}^{*}$ 0.016 0.028 0.052 0.012 0.024 0.056 0.008 0.044 0.080
$h=8$ $\mathscr{T}_{n}$ 0.028 0.064 0.092 0.028 0.056 0.092 0.016 0.044 0.094
$\mathscr{T}_{n}^{*}$ 0.004 0.036 0.064 0.012 0.040 0.064 0.008 0.020 0.048
7.1 $h=2$ $\mathscr{T}_{n}$ 0.036 0.128 0.164 0.064 0.116 0.164 0.040 0.112 0.132
$\mathscr{T}_{n}^{*}$ 0.004 0.028 0.076 0.016 0.040 0.072 0.024 0.048 0.092
$h=4$ $\mathscr{T}_{n}$ 0.020 0.076 0.092 0.036 0.088 0.116 0.028 0.088 0.116
$\mathscr{T}_{n}^{*}$ 0.008 0.028 0.048 0.012 0.040 0.084 0.000 0.036 0.088
$h=8$ $\mathscr{T}_{n}$ 0.024 0.052 0.076 0.012 0.044 0.072 0.024 0.048 0.060
$\mathscr{T}_{n}^{*}$ 0.012 0.048 0.090 0.008 0.050 0.086 0.000 0.020 0.072

### v2

@@ Appendix D Simulation Results
7.1 $h=2$ asy-p 0.037 0.071 0.099 0.033 0.061 0.095 0.048 0.085 0.124
chi-p 0.008 0.055 0.096 0.007 0.040 0.092 0.016 0.061 0.124
$h=4$ asy-p 0.046 0.083 0.109 0.030 0.064 0.097 0.046 0.078 0.118
chi-p 0.024 0.057 0.108 0.010 0.046 0.093 0.016 0.061 0.115
$h=8$ asy-p 0.079 0.111 0.138 0.031 0.059 0.079 0.026 0.047 0.082
chi-p 0.053 0.105 0.132 0.012 0.046 0.078 0.011 0.036 0.085
7.1 $h=2$ asy-p 0.022 0.050 0.079 0.024 0.056 0.084 0.034 0.075 0.102
chi-p 0.004 0.034 0.079 0.010 0.035 0.080 0.003 0.047 0.102
$h=4$ asy-p 0.040 0.071 0.104 0.027 0.057 0.086 0.026 0.066 0.104
chi-p 0.016 0.052 0.099 0.009 0.040 0.082 0.007 0.043 0.098
$h=8$ asy-p 0.069 0.102 0.130 0.025 0.049 0.072 0.017 0.042 0.089
chi-p 0.047 0.086 0.124 0.011 0.041 0.072 0.009 0.029 0.091
7.1 $h=2$ asy-p 0.029 0.069 0.087 0.035 0.070 0.104 0.028 0.072 0.106
chi-p 0.010 0.041 0.085 0.005 0.051 0.104 0.010 0.048 0.104
$h=4$ asy-p 0.039 0.073 0.103 0.035 0.066 0.102 0.034 0.067 0.107
chi-p 0.018 0.062 0.098 0.009 0.048 0.095 0.008 0.051 0.101
$h=8$ asy-p 0.069 0.113 0.146 0.022 0.053 0.083 0.019 0.046 0.092
chi-p 0.049 0.094 0.143 0.012 0.039 0.077 0.008 0.036 0.092
$d_{\lambda}=4$ $n=200$ $n=500$ $n=900$
@@
7.1 $h=2$ asy-p 0.036 0.064 0.085 0.036 0.062 0.088 0.040 0.074 0.097
chi-p 0.015 0.046 0.083 0.011 0.045 0.088 0.018 0.052 0.097
$h=4$ asy-p 0.046 0.089 0.116 0.037 0.060 0.092 0.040 0.074 0.104
chi-p 0.028 0.063 0.113 0.019 0.050 0.086 0.018 0.052 0.097
$h=8$ asy-p 0.097 0.139 0.170 0.042 0.061 0.083 0.024 0.054 0.086
chi-p 0.065 0.122 0.166 0.028 0.053 0.077 0.006 0.043 0.088
7.1 $h=2$ asy-p 0.028 0.053 0.082 0.026 0.055 0.075 0.032 0.071 0.099
chi-p 0.013 0.032 0.081 0.006 0.031 0.074 0.010 0.048 0.096
$h=4$ asy-p 0.044 0.079 0.109 0.027 0.054 0.073 0.032 0.063 0.095
chi-p 0.023 0.058 0.106 0.011 0.043 0.072 0.010 0.043 0.087
$h=8$ asy-p 0.070 0.110 0.140 0.028 0.054 0.073 0.026 0.053 0.088
chi-p 0.050 0.094 0.137 0.018 0.044 0.070 0.006 0.042 0.090
7.1 $h=2$ asy-p 0.029 0.052 0.070 0.027 0.061 0.086 0.027 0.054 0.077
chi-p 0.010 0.039 0.067 0.009 0.039 0.085 0.009 0.037 0.074
$h=4$ asy-p 0.055 0.093 0.128 0.027 0.058 0.089 0.031 0.060 0.089
chi-p 0.029 0.071 0.122 0.012 0.036 0.086 0.014 0.043 0.082
$h=8$ asy-p 0.090 0.128 0.157 0.029 0.053 0.075 0.022 0.044 0.085
chi-p 0.061 0.112 0.150 0.015 0.040 0.068 0.012 0.035 0.086
Table 6: Empirical sizes for $\mathbb{W}_{3}$, rejection probabilities at 1%, 5% and 10% levels.
$d_{\tau}=2$ $n=200$ $n=500$ $n=900$
@@
7.1 $h=2$ asy-p 0.036 0.064 0.085 0.036 0.062 0.088 0.040 0.074 0.097
chi-p 0.015 0.046 0.083 0.011 0.045 0.088 0.018 0.052 0.097
$h=4$ asy-p 0.046 0.089 0.116 0.037 0.060 0.092 0.040 0.074 0.104
chi-p 0.028 0.063 0.113 0.019 0.050 0.086 0.018 0.052 0.097
$h=8$ asy-p 0.097 0.139 0.170 0.042 0.061 0.083 0.024 0.054 0.086
chi-p 0.065 0.122 0.166 0.028 0.053 0.077 0.006 0.043 0.088
7.1 $h=2$ asy-p 0.028 0.053 0.082 0.026 0.055 0.075 0.032 0.071 0.099
chi-p 0.013 0.032 0.081 0.006 0.031 0.074 0.010 0.048 0.096
$h=4$ asy-p 0.044 0.079 0.109 0.027 0.054 0.073 0.032 0.063 0.095
chi-p 0.023 0.058 0.106 0.011 0.043 0.072 0.010 0.043 0.087
$h=8$ asy-p 0.070 0.110 0.140 0.028 0.054 0.073 0.026 0.053 0.088
chi-p 0.050 0.094 0.137 0.018 0.044 0.070 0.006 0.042 0.090
7.1 $h=2$ asy-p 0.029 0.052 0.070 0.027 0.061 0.086 0.027 0.054 0.077
chi-p 0.010 0.039 0.067 0.009 0.039 0.085 0.009 0.037 0.074
$h=4$ asy-p 0.055 0.093 0.128 0.027 0.058 0.089 0.031 0.060 0.089
chi-p 0.029 0.071 0.122 0.012 0.036 0.086 0.014 0.043 0.082
$h=8$ asy-p 0.090 0.128 0.157 0.029 0.053 0.075 0.022 0.044 0.085
chi-p 0.061 0.112 0.150 0.015 0.040 0.068 0.012 0.035 0.086
Table 7: Empirical sizes for $\mathbb{W}_{4}$, rejection prob. at 1%, 5% and 10% levels. Testing $\delta(z)$.
$d_{\lambda}=1$ $n=200$ $n=500$ $n=900$
@@
7.1 $h=2$ asy-p 0.065 0.080 0.129 0.053 0.086 0.110 0.050 0.072 0.090
chi-p 0.022 0.053 0.124 0.027 0.069 0.109 0.021 0.059 0.089
$h=4$ asy-p 0.042 0.085 0.112 0.071 0.106 0.130 0.047 0.072 0.098
chi-p 0.039 0.071 0.103 0.026 0.090 0.126 0.029 0.061 0.092
$h=8$ asy-p 0.169 0.124 0.159 0.019 0.060 0.074 0.022 0.039 0.071
chi-p 0.144 0.101 0.154 0.025 0.054 0.071 0.016 0.036 0.080
7.1 $h=2$ asy-p 0.076 0.084 0.144 0.056 0.078 0.090 0.041 0.060 0.070
chi-p 0.031 0.055 0.137 0.028 0.071 0.087 0.017 0.053 0.069
$h=4$ asy-p 0.134 0.085 0.133 0.075 0.097 0.115 0.054 0.077 0.093
chi-p 0.099 0.064 0.125 0.015 0.065 0.111 0.014 0.060 0.091
$h=8$ asy-p 0.174 0.124 0.153 0.053 0.081 0.094 0.025 0.045 0.080
chi-p 0.142 0.103 0.150 0.022 0.072 0.092 0.015 0.036 0.083
7.1 $h=2$ asy-p 0.032 0.059 0.126 0.058 0.080 0.098 0.051 0.077 0.095
chi-p 0.018 0.039 0.122 0.033 0.067 0.094 0.017 0.062 0.093
$h=4$ asy-p 0.072 0.094 0.141 0.060 0.100 0.129 0.055 0.084 0.104
chi-p 0.095 0.070 0.137 0.041 0.077 0.122 0.016 0.070 0.101
$h=8$ asy-p 0.063 0.089 0.151 0.047 0.067 0.083 0.025 0.049 0.076
chi-p 0.109 0.070 0.147 0.020 0.063 0.078 0.016 0.036 0.088
$d_{\lambda}=2$ $n=200$ $n=500$ $n=900$
@@
7.1 $h=2$ asy-p 0.086 0.117 0.135 0.061 0.086 0.099 0.028 0.053 0.076
chi-p 0.061 0.101 0.132 0.036 0.075 0.098 0.016 0.041 0.082
$h=4$ asy-p 0.122 0.163 0.187 0.063 0.084 0.111 0.041 0.059 0.068
chi-p 0.091 0.144 0.184 0.044 0.072 0.107 0.030 0.048 0.082
$h=8$ asy-p 0.145 0.195 0.227 0.043 0.065 0.089 0.028 0.047 0.058
chi-p 0.109 0.177 0.225 0.030 0.057 0.083 0.018 0.038 0.081
7.1 $h=2$ asy-p 0.083 0.113 0.141 0.057 0.081 0.098 0.034 0.051 0.067
chi-p 0.052 0.101 0.136 0.036 0.067 0.096 0.018 0.042 0.085
$h=4$ asy-p 0.117 0.154 0.183 0.060 0.090 0.113 0.028 0.046 0.063
chi-p 0.087 0.138 0.176 0.045 0.078 0.109 0.019 0.040 0.081
$h=8$ asy-p 0.134 0.188 0.218 0.041 0.067 0.086 0.027 0.046 0.068
chi-p 0.105 0.167 0.214 0.032 0.060 0.085 0.016 0.036 0.081
7.1 $h=2$ asy-p 0.093 0.131 0.148 0.045 0.066 0.091 0.036 0.058 0.069
chi-p 0.057 0.114 0.145 0.030 0.056 0.087 0.020 0.048 0.087
$h=4$ asy-p 0.135 0.177 0.203 0.064 0.088 0.107 0.030 0.047 0.076
chi-p 0.090 0.163 0.200 0.043 0.074 0.104 0.018 0.038 0.084
$h=8$ asy-p 0.148 0.202 0.228 0.060 0.080 0.092 0.022 0.036 0.072
chi-p 0.115 0.179 0.227 0.039 0.077 0.089 0.016 0.030 0.082